4H20 + 3Fe -> Fe3O4 + 4H2
What volume of H2 at a pressure of 1.12 atm and a temperature of 40C can be prepared from the reaction of 20g of H2O
1
Expert's answer
2021-10-10T10:16:50-0400

Pressure = 1.12atm

Temperature = 40°C + 273 = 313K

"20g H_2O ×\frac{1 mole H_2}{1 mole H_2O}=1.1099molH_2"


"V=\frac{nRT}{P}"


"V=(1.1099)(\frac{0.112Latm}{mol×K})(313K)÷112atm =34.74LH_2"